# Stochastic Optimization with Random Search

### v1

@@ Appendix B Analysis of the General Algorithm
Let $f$ satisfy equation 13 and let the random directions ${\bm{s}}_{t}\sim{\mathcal{D}}$ satisfy Assumption 2, namely $\mathbb{E}\|{\bm{s}}_{t}\|^{2}=1$ and $\mathbb{E}[\,|\langle\nabla f({\bm{x}}),{\bm{s}}_{t}\rangle|\,]\geq\mu_{{\mathcal{D}}}\|\nabla f({\bm{x}})\|.$ Then, for any step size $\eta_{t}\leq\mu_{{\mathcal{D}}}/L_{1}$, the update defined above satisfies
$\mathbb{E}[f({\bm{x}}_{t+1})]$ $\leq\mathbb{E}[f({\bm{x}}_{t})]-\tfrac{\mu_{{\mathcal{D}}}}{2}\,\eta_{t}\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|+\tfrac{L_{0}}{2}\,\eta_{t}^{2}+2\,\mathbb{E}\!\left[\big|M_{t}^{+}-M_{t}^{-}-(f({\bm{x}}_{t}^{+})-f({\bm{x}}_{t}^{-}))\big|\right],$ (15)
where ${\bm{x}}_{t}^{\pm}={\bm{x}}_{t}\pm\eta_{t}{\bm{s}}_{t}$.
@@ C.2 Analysis under individual Smoothness
For $\mathcal{P}$-a.e. $\xi$, the mapping $x\mapsto f_{\xi}({\bm{x}})$ is $(L_{0}(\xi),L_{1}(\xi))$-smooth.
Assumption 6.
The random smoothness parameters satisfy
$\mathcal{L}_{0}:=\sqrt{\mathbb{E}_{\xi}\!\big[L_{0}(\xi)^{2}\big]}<\infty,\qquad L_{1}(\xi)\leq\mathcal{L}_{1}<\infty\ \text{for $\mathcal{P}$-a.e.\ $\xi$}.$
Assumption 7.
There exists $\sigma<\infty$ such that for all $x\in\mathbb{R}^{d}$,
@@
To simplify the analysis, we assume that $\mathcal{D}$ is the uniform distribution over the unit sphere of $\mathbb{R}^{d}$.
@@
Let $u\in\mathbb{S}^{d-1}$ be uniformly distributed on the unit sphere and let $v\in\mathbb{R}^{d}$. Then
$\mathbb{E}\big[\langle v,u\rangle^{2}\big]\;=\;\frac{\|v\|^{2}}{d}.$
@@
Let ${\bm{x}}_{t}^{\pm}={\bm{x}}_{t}\pm\eta_{t}s_{t}$ with $\eta_{t}\leq 1/{\mathcal{L}}_{1}$ and assume that assumptions 5,6 and 7 hold. With the minibatch estimators $M_{t}^{\pm}=\frac{1}{b}\sum_{j=1}^{b}f_{\xi^{t}_{j}}({\bm{x}}_{t}^{\pm})$, define
$\Delta_{t}\;:=\;(M_{t}^{+}-M_{t}^{-})-\big(f({\bm{x}}_{t}^{+})-f({\bm{x}}_{t}^{-})\big).$
Then, conditioning only on ${\bm{x}}_{t}$ and taking expectation over both the minibatch and the direction, we have
$\mathbb{E}\big[\,|\Delta_{t}|\,\big|\,{\bm{x}}_{t}\big]\;\leq\;\underbrace{\frac{2\eta_{t}}{\sqrt{db}}\,\sigma}_{\textstyle\text{variance term}}\;+\;\underbrace{\frac{4\sqrt{2}\,\eta_{t}^{2}}{\sqrt{b}}\,\big({\mathcal{L}}_{0}+{\mathcal{L}}_{1}\sigma+{\mathcal{L}}_{1}\|\nabla f({\bm{x}}_{t})\|\big)}_{\textstyle\text{bias/smoothness term}}.$
@@
Let $y_{t}={\bm{x}}_{t}^{+}$, $z_{t}={\bm{x}}_{t}^{-}$, $\delta_{t}=y_{t}-z_{t}=2\eta_{t}s_{t}$, and $u_{t}=\delta_{t}/\|\delta_{t}\|=s_{t}/\|s_{t}\|$. For each $j$, by the fundamental theorem of calculus,
$f_{\xi^{t}_{j}}(y_{t})-f_{\xi^{t}_{j}}(z_{t})=\|\delta_{t}\|\int_{0}^{1}\big\langle\nabla f_{\xi^{t}_{j}}(z_{t}+x\delta_{t}),\,u_{t}\big\rangle\,dx.$
Averaging over $j=1,\dots,b$ and subtracting the mean gives
$\Delta_{t}=\frac{\|\delta_{t}\|}{b}\sum_{j=1}^{b}\Big(D^{(t)}_{j}-\mathbb{E}_{\xi}[D^{(t)}_{j}]\Big),\qquad D^{(t)}_{j}:=\int_{0}^{1}\big\langle\nabla f_{\xi^{t}_{j}}(z_{t}+x\delta_{t}),\,u_{t}\big\rangle\,dx.$
Decompose $D^{(t)}_{j}=A^{(t)}_{j}+R^{(t)}_{j}$ with
$A^{(t)}_{j}:=\big\langle\nabla f_{\xi^{t}_{j}}(z_{t}),\,u_{t}\big\rangle,\qquad R^{(t)}_{j}:=\int_{0}^{1}\big\langle\nabla f_{\xi^{t}_{j}}(z_{t}+x\delta_{t})-\nabla f_{\xi^{t}_{j}}(z_{t}),\,u_{t}\big\rangle\,dx.$
Hence
@@
We bound the two contributions separately.
The variance term $(\star)$.
Conditioning on ${\bm{x}}_{t}$ and using independence of $u_{t}$ and $\{\xi^{t}_{j}\}$,
$\mathbb{E}\!\left[\left|\frac{1}{b}\sum_{j=1}^{b}\big(A^{(t)}_{j}-\mathbb{E}[A^{(t)}_{j}]\big)\right|\,\Big|\,{\bm{x}}_{t}\right]$ $\leq\sqrt{\mathrm{Var}\!\left(\frac{1}{b}\sum_{j=1}^{b}A^{(t)}_{j}\,\Big|\,{\bm{x}}_{t}\right)}=\frac{1}{\sqrt{b}}\sqrt{\mathrm{Var}\!\left(A^{(t)}_{1}\,\Big|\,{\bm{x}}_{t}\right)}$
@@
$=\frac{1}{\sqrt{b}}\sqrt{\mathbb{E}_{u_{t},\xi}\Big[\langle\nabla f_{\xi}(z_{t})-\nabla f(z_{t}),u_{t}\rangle^{2}\Big]}$
$=\frac{1}{\sqrt{b}}\sqrt{\mathbb{E}_{\xi}\!\left[\frac{\|\nabla f_{\xi}(z_{t})-\nabla f(z_{t})\|^{2}}{d}\right]}\;\leq\;\frac{\sigma}{\sqrt{db}},$
where we used Lemma C.1 to pass from the inner product to $\|\,\cdot\,\|^{2}/d$.
Multiplying by $\|\delta_{t}\|=2\eta_{t}\|s_{t}\|$ and using $\|s_{t}\|=1$ yields the first claimed term $\frac{2\eta_{t}}{\sqrt{db}}\,\sigma$.
The smoothness term $(\diamond)$.
Using the $(L_{0}(\xi),L_{1}(\xi))$-smoothness of $f_{\xi}$ to get, for $x\in[0,1]$,
$\big\|\nabla f_{\xi}(z_{t}+x\delta_{t})-\nabla f_{\xi}(z_{t})\big\|\leq\big(L_{0}(\xi)+L_{1}(\xi)\|\nabla f_{\xi}(z_{t})\|\big)\,x\,\|\delta_{t}\|.$
Using Assumption 5, for all $x\in(0,1)$ and for $\mathbb{P}$-almost every $\xi$,
$\bigl\|\nabla f_{\xi}(z_{t}+x\delta_{t})-\nabla f_{\xi}(z_{t})\bigr\|_{2}\;\leq\;L_{0}(\xi)\,x\|\delta_{t}\|_{2}\;+\;L_{1}(\xi)\,\|\nabla f_{\xi}(z_{t})\|_{2}\,x\,\|\delta_{t}\|_{2}.$
Then,
$\bigl|R_{1}^{(t)}\bigr|$ $\leq\int_{0}^{1}\!\Bigl(L_{0}(\xi_{1}^{t})\,x\|\delta_{t}\|_{2}+L_{1}(\xi_{1}^{t})\,\|\nabla f_{\xi_{1}^{t}}(z_{t})\|_{2}\,x\,\|\delta_{t}\|_{2}\Bigr)\,dx$
$=\frac{L_{0}(\xi_{1}^{t})\,\|\delta_{t}\|_{2}}{2}\;+\;\frac{L_{1}(\xi_{1}^{t})\,\|\nabla f_{\xi_{1}^{t}}(z_{t})\|_{2}\,\|\delta_{t}\|_{2}}{2}.$
Using the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$, we obtain
$\mathbb{E}\!\left[\bigl(R_{1}^{(t)}\bigr)^{2}\,\middle|\,x_{t},s_{t}\right]\;\leq\;\frac{\|\delta_{t}\|_{2}^{2}}{2}\,\mathbb{E}_{\xi}\!\bigl[L_{0}(\xi)^{2}\bigr]\;+\;\frac{1}{2}\,\mathbb{E}_{\xi}\!\bigl[L_{1}(\xi)^{2}\,\|\nabla f_{\xi}(z_{t})\|_{2}^{2}\bigr]\,\|\delta_{t}\|_{2}^{2}.$
Using Assumptions 6 and 7, we further obtain
$\mathbb{E}\!\left[\bigl(R_{1}^{(t)}\bigr)^{2}\,\middle|\,x_{t},s_{t}\right]\;\leq\;\frac{(\mathcal{L}_{0})^{2}\,\|\delta_{t}\|_{2}^{2}}{2}\;+\;\frac{(\mathcal{L}_{1})^{2}\,\|\delta_{t}\|_{2}^{2}}{2}\,\bigl(\sigma^{2}+\|\nabla f(z_{t})\|_{2}^{2}\bigr).$
Therefore, using Jensen’s inequality:
$\mathbb{E}\!\left[\left|\frac{1}{b}\sum_{j=1}^{b}\!\bigl(R_{j}^{(t)}-\mathbb{E}[R_{j}^{(t)}]\bigr)\right|\,\middle|\,x_{t},s_{t}\right]$ $\leq\sqrt{\operatorname{Var}_{\xi_{1},\ldots,\xi_{b}}\!\left(\frac{1}{b}\sum_{j=1}^{b}\!R_{j}^{(t)}\,\middle|\,x_{t},s_{t}\right)}$
$\leq\frac{1}{\sqrt{b}}\,\sqrt{\mathbb{E}\!\left[(R_{1}^{(t)})^{2}\,\middle|\,x_{t},s_{t}\right]}$
$\leq\frac{1}{\sqrt{2b}}\Bigl(\mathcal{L}_{0}\,\|\delta_{t}\|+\mathcal{L}_{1}\,\|\delta_{t}\|\,(\sigma+\|\nabla f(z_{t})\|)\Bigr)$
$\leq\frac{\sqrt{2}\eta_{t}}{\sqrt{b}}\Bigl(\mathcal{L}_{0}+\mathcal{L}_{1}(\sigma+\|\nabla f(z_{t})\|)\Bigr).$
We deduce that:
$\mathbb{E}_{\xi_{1},\ldots,\xi_{b}}\!\Big[\big|M_{t}^{+}-M_{t}^{-}-(f(y_{t})-f(z_{t}))\big|\,\Big|\,x_{t}\Big]$ $\leq\frac{2\eta_{t}}{\sqrt{db}}\,\sigma\;+\;\frac{2\sqrt{2}\eta_{t}^{2}}{\sqrt{b}}\Bigl(\mathcal{L}_{0}+\mathcal{L}_{1}(\sigma+\|\nabla f(z_{t})\|)\Bigr).$
Using Assumption 5, for $\mathbb{P}$-almost every $\xi$, we have:
$\|\nabla f_{\xi}(z_{t})-\nabla f_{\xi}(x_{t})\|\;\leq\;\big(L_{0}(\xi)+L_{1}(\xi)\|\nabla f_{\xi}(x_{t})\|\big)\,\eta_{t}\|s_{t}\|.$
Using Jensen’s inequality, assumptions 6 and 7, we obtain:
$\|\nabla f(z_{t})-\nabla f(x_{t})\|$ $=\big\|\mathbb{E}_{\xi}[\nabla f_{\xi}(z_{t})-\nabla f_{\xi}(x_{t})]\big\|$
$\leq\mathbb{E}_{\xi}\big[\|\nabla f_{\xi}(z_{t})-\nabla f_{\xi}(x_{t})\|\big]$
$\leq\mathbb{E}_{\xi}\!\Big[(L_{0}(\xi)+L_{1}(\xi)\|\nabla f_{\xi}(x_{t})\|)\,\eta_{t}\Big]$
$\leq\eta_{t}\!\left(\mathcal{L}_{0}+\mathcal{L}_{1}\,\mathbb{E}_{\xi}\|\nabla f_{\xi}(x_{t})\|\right)$
$\leq\eta_{t}\!\left(\mathcal{L}_{0}+\mathcal{L}_{1}\sqrt{\mathbb{E}_{\xi}\|\nabla f_{\xi}(x_{t})\|^{2}}\right)$
$\leq\eta_{t}\!\left(\mathcal{L}_{0}+\mathcal{L}_{1}\sqrt{\sigma^{2}+\|\nabla f(x_{t})\|^{2}}\right)$
$\leq\eta_{t}\!\left(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma+\mathcal{L}_{1}\|\nabla f(x_{t})\|\right).$
If $\eta_{t}\leq\tfrac{1}{\mathcal{L}_{1}}$, then by the triangle inequality:
$\|\nabla f(z_{t})\|$ $\leq\|\nabla f(x_{t})\|+\|\nabla f(z_{t})-\nabla f(x_{t})\|$
$\leq 2\|\nabla f(x_{t})\|+\frac{\mathcal{L}_{0}+\mathcal{L}_{1}\sigma}{\mathcal{L}_{1}}.$
We conclude that:
$\mathbb{E}\!\left[\,|\Delta_{t}|\,\middle|\,x_{t}\right]\;\leq\;\frac{2\eta_{t}}{\sqrt{db}}\,\sigma\;+\;\frac{4\sqrt{2}\,\eta_{t}^{\,2}}{\sqrt{b}}\,\bigl(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma+\mathcal{L}_{1}\|\nabla f(x_{t})\|\bigr).$
@@
Assume that assumptions 2, 5, 6, and 7 hold and the step size $\eta_{t}=\eta\leq\;\min\!\left\{\frac{\mu_{\mathcal{D}}}{\mathcal{L}_{1}},\,\frac{\mu_{\mathcal{D}}\sqrt{b}}{32\sqrt{2}\,\mathcal{L}_{1}}\right\}$. We have for all $t\geq 0$:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|\;\leq\;\frac{4F_{0}}{\mu_{\mathcal{D}}\eta\,T}\;+\;\frac{2\mathcal{L}_{0}+2\mathcal{L}_{1}\sigma}{\mu_{\mathcal{D}}}\,\eta\;+\;\frac{16}{\mu_{\mathcal{D}}}\,\frac{\sigma}{\sqrt{db}}\;+\;\frac{32\sqrt{2}}{\mu_{\mathcal{D}}}\,\frac{\eta}{\sqrt{b}}\,(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma),$
where $F_{0}:=f({\bm{x}}_{0})-f^{\star}.$
@@
Let $t\geq 0$. We have for all $x,y\in\mathbb{R}^{d}:$
$\|\nabla f(y)-\nabla f(x)\|$ $=\|\,\mathbb{E}_{\xi}[\nabla f_{\xi}(y)-\nabla f_{\xi}(x)]\,\|$
$\leq\mathbb{E}_{\xi}[\,\|\nabla f_{\xi}(y)-\nabla f_{\xi}(x)\|\,]$
$\leq\mathbb{E}_{\xi}[L_{0}(\xi)+L_{1}(\xi)\|\nabla f_{\xi}(x)\|]\,\|x-y\|$
$\leq(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma+\mathcal{L}_{1}\|\nabla f(x)\|)\|x-y\|.$
Then $f$ is $(\mathcal{L}_{0}+\mathcal{L}_{1}\sigma,\mathcal{L}_{1})$ smooth, and by applying theorem B.1 , since $\eta_{t}=\eta\leq\mu_{\mathcal{D}}/\mathcal{L}_{1}$, we have:
$\mathbb{E}[f({\bm{x}}_{t+1})]\leq\mathbb{E}[f({\bm{x}}_{t})]-\tfrac{\mu_{{\mathcal{D}}}}{2}\,\eta\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|+\tfrac{\mathcal{L}_{0}+\mathcal{L}_{1}\sigma}{2}\,\eta^{2}+2\,\mathbb{E}\!\left[\big|M_{t}^{+}-M_{t}^{-}-(f({\bm{x}}_{t}^{+})-f({\bm{x}}_{t}^{-}))\big|\right].$

### v2

@@ Appendix B Analysis of the General Algorithm
Let $f$ satisfy equation 13 and let the random directions ${\bm{s}}_{t}\sim{\mathcal{D}}$ satisfy Assumption 2, namely $\mathbb{E}\|{\bm{s}}_{t}\|^{2}=1$ and $\mathbb{E}[\,|\langle\nabla f({\bm{x}}),{\bm{s}}_{t}\rangle|\,]\geq\mu_{{\mathcal{D}}}\|\nabla f({\bm{x}})\|.$ Then, for any step size $\eta_{t}\leq\mu_{{\mathcal{D}}}/L_{1}$, the update defined above satisfies $\mathbb{E}[f({\bm{x}}_{t+1})]$ $\leq\mathbb{E}[f({\bm{x}}_{t})]-\tfrac{\mu_{{\mathcal{D}}}}{2}\,\eta_{t}\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|+\tfrac{L_{0}}{2}\,\eta_{t}^{2}+2\,\mathbb{E}\!\left[\big|M_{t}^{+}-M_{t}^{-}-(f({\bm{x}}_{t}^{+})-f({\bm{x}}_{t}^{-}))\big|\right],$ (15) where ${\bm{x}}_{t}^{\pm}={\bm{x}}_{t}\pm\eta_{t}{\bm{s}}_{t}$.
@@ C.2 Analysis under individual Smoothness
For $\mathcal{P}$-a.e. $\xi$, the mapping $x\mapsto f_{\xi}({\bm{x}})$ is $(L_{0},L_{1})$-smooth.
Assumption 6 (Bounded gradient variance).
There exists $\sigma\geq 0$ such that for all ${\bm{x}}\in\mathbb{R}^{d}$,
@@
Lemma C.1 (Equivalent second-moment form).
Assumption 6 holds if and only if, for all ${\bm{x}}\in\mathbb{R}^{d}$, $\mathbb{E}_{\xi}\!\big[\|\nabla f_{\xi}({\bm{x}})\|^{2}\big]\ \leq\ \|\nabla f({\bm{x}})\|^{2}+\sigma^{2}.$
Proof.
Let $G_{\xi}({\bm{x}}):=\nabla f_{\xi}({\bm{x}})$ and $G({\bm{x}}):=\nabla f({\bm{x}})=\mathbb{E}_{\xi}[G_{\xi}({\bm{x}})]$. The variance decomposition identity gives
$\mathbb{E}_{\xi}\!\big[\|G_{\xi}({\bm{x}})\|^{2}\big]\;=\;\|G({\bm{x}})\|^{2}+\mathbb{E}_{\xi}\!\big[\|G_{\xi}({\bm{x}})-G({\bm{x}})\|^{2}\big].$
Therefore
$\mathbb{E}_{\xi}\!\big[\|G_{\xi}({\bm{x}})-G({\bm{x}})\|^{2}\big]\leq\sigma^{2}\quad\Longleftrightarrow\quad\mathbb{E}_{\xi}\!\big[\|G_{\xi}({\bm{x}})\|^{2}\big]\leq\|G({\bm{x}})\|^{2}+\sigma^{2}.$
@@
By sampling independent samples $\xi_{1},\ldots,\xi_{b}$ from $\mathcal{P},$ we have for all $x,y\in\mathbb{R}^{d}:$ $\mathbb{E}\lVert\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(y)-\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(x)\rVert\leq\big(L_{0}+L_{1}\sigma+L_{1}\|\nabla f(x)\|\big)\,\|x-y\|,$ and we have also: $\lVert\nabla f(y)-\nabla f(x)\rVert\leq\big(L_{0}+L_{1}\sigma+L_{1}\|\nabla f(x)\|\big)\,\|x-y\|.$
Proof.
Let $x,y\in\mathbb{R}^{d}$. Using Assumption 5, for $\mathbb{P}$-almost every $\xi$, we have:
$\|\nabla f_{\xi}(y)-\nabla f_{\xi}(x)\|\;\leq\;\big(L_{0}+L_{1}\|\nabla f_{\xi}(x)\|\big)\,\|y-x\|.$
By sampling independent samples $\xi_{1},\ldots,\xi_{b}$ from $\mathcal{P},$ we have:
$\lVert\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(y)-\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(x)\rVert$ $\leq\frac{\sum_{j=1}^{b}\lVert\nabla f_{\xi_{j}}(y)-\nabla f_{\xi_{j}}(x)\rVert}{b}$
$\leq\bigg(L_{0}+L_{1}\frac{\sum_{j=1}\|\nabla f_{\xi_{j}}(x)\|}{b}\bigg)\|x-y\|.$
This implies that:
$\mathbb{E}\lVert\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(y)-\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(x)\rVert$ $\leq\big(L_{0}+L_{1}\mathbb{E}\|\nabla f_{\xi_{1}}(x)\|\big)\|x-y\|$
$\leq L_{0}+L_{1}\sqrt{\mathbb{E}\big(\|\nabla f_{\xi_{1}}(x)\|^{2}}\big).\,\|x-y\|$
$\leq\big(L_{0}+L_{1}\sqrt{\sigma^{2}+\|\nabla f(x)\|^{2}}\big)\|x-y\|$
$\leq\big(L_{0}+L_{1}\sigma+L_{1}\|\nabla f(x)\|\big)\|x-y\|.$
For the second claim, using Jensen’s inequality and assumption 6, we obtain:
$\|\nabla f(y)-\nabla f(x)\|$ $=\big\|\mathbb{E}_{\xi}[\nabla f_{\xi}(y)-\nabla f_{\xi}(x)]\big\|$
$\leq\mathbb{E}_{\xi}\big[\|\nabla f_{\xi}(y)-\nabla f_{\xi}(x)\|\big]$
$\leq\mathbb{E}_{\xi}\!\Big[L_{0}+L_{1}\|\nabla f_{\xi}(x)\|\ \,\Big]\|x-y\|$
$\leq\Big[L_{0}+L_{1}\mathbb{E}\|\nabla f_{\xi}(x)\|\Big]\|x-y\|$
$\leq\!\left(L_{0}+L_{1}\sqrt{\sigma^{2}+\|\nabla f(x)\|^{2}}\right)\|x-y\|$
$\leq\!\left(L_{0}+L_{1}\sigma+L_{1}\|\nabla f(x)\|\right)\|x-y\|.$
For simplicity, we assume throughout the rest of this section that the search distribution $\mathcal{D}$ is uniform on the unit sphere.
Lemma C.3.
Let $u\in\mathbb{S}^{d-1}$ be uniformly distributed on the unit sphere and let $v\in\mathbb{R}^{d}$. Then $\mathbb{E}\big[\langle v,u\rangle^{2}\big]\;=\;\frac{\|v\|^{2}}{d}.$
@@
Let ${\bm{x}}_{t}^{\pm}:={\bm{x}}_{t}\pm\eta s_{t}$, and assume Assumptions 5 and 6 hold. Suppose the search direction $s_{t}$ is drawn uniformly from the unit sphere $\mathbb{S}^{d-1}$. Let $\xi_{t,1},\ldots,\xi_{t,b}$ be i.i.d. samples from $\mathcal{P}$, and define the mini-batch estimators $M_{t}^{\pm}\;:=\;\frac{1}{b}\sum_{j=1}^{b}f_{\xi_{t,j}}({\bm{x}}_{t}^{\pm}).$ Then $\mathbb{E}\big[\big|(M^{+}-M^{-})-(f(x_{t}^{+})-f(x_{t}^{-}))\big|\,\big|x_{t}\big]\leq\frac{2\eta\sigma}{\sqrt{db}}+2\eta^{2}\big(L_{0}+L_{1}\sigma+L_{1}\|\nabla f(x_{t})\|\big).$
@@
Denote: $x_{t}^{\pm}=x_{t}\pm\eta s_{t}$. By the fundamental theorem of calculus, we have:
$\begin{cases}f_{\xi}(x_{t}^{+})-f_{\xi}(x_{t})\;=\;\int_{0}^{1}\!\!\!\big\langle\nabla f_{\xi}(x_{t}+u\eta s_{t}),\,\eta s_{t}\big\rangle\,du,\\ f_{\xi}(x_{t}^{-})-f_{\xi}(x_{t})\;=\;\int_{0}^{-1}\!\!\!\big\langle\nabla f_{\xi}(x_{t}+u\eta s_{t}),\,\eta s_{t}\big\rangle\,du.\end{cases}\Longrightarrow f_{\xi}(x_{t}^{+})-f_{\xi}(x_{t}^{-})=\int_{-1}^{1}\!\!\!\big\langle\nabla f_{\xi}(x_{t}+u\eta s_{t}),\,\eta s_{t}\big\rangle\,du.$
Averaging over the minibatch and subtracting the population identity gives:
$(M^{+}-M^{-})-(f(x_{t}^{+})-f(x_{t}^{-}))$ $=\eta\int_{-1}^{1}\!\Big\langle\underbrace{\tfrac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(x_{t}+u\eta s_{t})-\nabla f(x_{t}+u\eta s_{t})}_{=:~\Delta(x_{t},s_{t})},\,s_{t}\Big\rangle\,du.$ ($\star$)
Denote: $\overline{g}(x):=\frac{1}{b}\sum_{j=1}^{b}\nabla f_{\xi_{j}}(x)$. We have:
$\Delta(x_{t},s_{t})=\big(\overline{g}(x_{t})-\nabla f(x_{t})\big)+\big(\overline{g}(x_{t}+u\eta s_{t})-\overline{g}(x_{t})\big)-\big(\nabla f(x_{t}+u\eta s_{t})-\nabla f(x_{t})\big).$
Plugging this into $(\star)$ and using the triangle inequality yields
$\big|(M^{+}-M^{-})-(f(x_{t}^{+})-f(x_{t}^{-}))\big|$ $\leq 2\eta\,\big|\langle\overline{g}(x_{t})-\nabla f(x_{t}),\,s_{t}\rangle\big|$
$\qquad+\eta\!\int_{-1}^{1}\!\Big(\big|\langle\overline{g}(x_{t}+u\eta s_{t})-\overline{g}(x_{t}),s_{t}\rangle\big|$
$\qquad+\big|\langle\nabla f(x_{t}+u\eta s_{t})-\nabla f(x_{t}),s_{t}\rangle\big|\Big)\,du.$ ($\dagger$)
By denoting $y_{u,t}=x_{t}+u\eta s_{t}.$ By lemma C.2, we have:
@@
Using Cauchy–Schwarz inequality and lemma C.3 we have:
$\mathbb{E}[\big|\langle\overline{g}(x_{t})-\nabla f(x_{t}),\,s_{t}\rangle\big|\,|x_{t},\xi_{1},\ldots,\xi_{b}]$ $\leq\sqrt{\mathbb{E}[\langle\overline{g}(x_{t})-\nabla f(x_{t}),\,s_{t}\rangle^{2}\,|x_{t},\xi_{1},\ldots,\xi_{b}]}$
$=\frac{\|\overline{g}(x_{t})-\nabla f(x_{t})\|}{\sqrt{d}}$
We have also: $\mathbb{E}\big[\|\overline{g}(x_{t})-\nabla f(x_{t})\|\,|x_{t}\big]\leq\sqrt{\mathbb{E}\big[\|\overline{g}(x_{t})-\nabla f(x_{t})\|^{2}\,|x_{t}\big]}=\sqrt{\frac{1}{b}\mathbb{E}\big[\|\nabla f_{\xi_{1}}(x_{t})-\nabla f(x_{t})\|^{2}\,|x_{t}\big]}.$ This implies that: $\mathbb{E}\big[\|\overline{g}(x_{t})-\nabla f(x_{t})\|\,|x_{t}\big]\leq\frac{\sigma}{\sqrt{b}}.$ It follows that: $\mathbb{E}\big[\frac{\|\overline{g}(x_{t})-\nabla f(x_{t})\|}{\sqrt{d}}\,|x_{t}\big]\leq\frac{\sigma}{\sqrt{db}}.$ Therefore:
@@
$\mathbb{E}\big[\big|\langle\overline{g}(x_{t})-\nabla f(x_{t}),\,s_{t}\rangle\big|\,|x_{t}\big]\leq\frac{\sigma}{\sqrt{db}}.$
Combining the bounds in 25 with ($\dagger$) and the inequality above, we obtain:
$\mathbb{E}\big[\big|(M^{+}-M^{-})-(f(x_{t}^{+})-f(x_{t}^{-}))\big|\,\big|x_{t}\big]\leq\frac{2\eta\sigma}{\sqrt{db}}+2\eta^{2}\big(L_{0}+L_{1}\sigma+L_{1}\|\nabla f(x_{t})\|\big).$
@@
Assume that assumptions 5, and 6 hold, the search directions sampled from the uniform distribution over the unit sphere, and the step size $\eta\leq\;\frac{\mu_{\mathcal{D}}}{16L_{1}}$. We have for all $t\geq 0$: $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla f({\bm{x}}_{t})\big\|\;\leq\;\frac{4F_{0}}{\mu_{\mathcal{D}}\eta\,T}\;+\;\frac{16}{\mu_{\mathcal{D}}}\,\frac{\sigma}{\sqrt{db}}\;+\;\frac{18\eta}{\mu_{\mathcal{D}}}\,(L_{0}+L_{1}\sigma).$ where $F_{0}:=f({\bm{x}}_{0})-f^{\star}.$
@@
Let $t\geq 0$. Using lemma C.2, $f$ is $(L_{0}+L_{1}\sigma,L_{1})$ smooth, and by applying theorem B.1 , since $\eta\leq\mu_{\mathcal{D}}/L_{1}$, we have:
$\mathbb{E}[f({\bm{x}}_{t+1})]\leq\mathbb{E}[f({\bm{x}}_{t})]-\tfrac{\mu_{{\mathcal{D}}}}{2}\,\eta\,\mathbb{E}\|\nabla f({\bm{x}}_{t})\|+\tfrac{L_{0}+L_{1}\sigma}{2}\,\eta^{2}+2\,\mathbb{E}\!\left[\big|M_{t}^{+}-M_{t}^{-}-(f({\bm{x}}_{t}^{+})-f({\bm{x}}_{t}^{-}))\big|\right].$